### v2

# Merging RLBWTs adaptively

Abstract
We show how we can merge two run-length compressed Burrows-Wheeler Transforms (RLBWTs) into a run-length compressed extended Burrows-Wheeler Transform (eBWT) in $O(r)$ space and $\tilde{O}(r+L)$ time, where $r$ is the number of runs in the final eBWT and $L$ is the sum of the longest common prefix (LCP) values at the beginnings of those runs.
1 Introduction
The Burrows-Wheeler Transform (BWT) [3] is an important tool in modern genomics. As DNA sequencing technologies have advanced and researchers have started working with pangenomic references, they have started using run-length compressed BWTs (RLBWTs). There are good algorithms for building RLBWTs for massive pangenomic references in which all the genomes are similar [2, 5] but the case when the reference contains many genomes from each of many species is still challenging. One possible solution is to build a separate RLBWT for each species and then merge them into a single RLBWT or run-length compressed extended BWT (eBWT) [6].
As far as we know, Oliva et al. [7] were the first to study merging RLBWTs, building on Bannai, Gagie and I’s [1] work on dynamic r-indexes. Their merging algorithm uses compressed space but time at least linear in the size of the uncompressed inputs. Recently, Díaz-Domínguez et al. [4] gave an algorithm for merging BWTs that could be extended to RLBWTs and works fairly well in practice but has no good worst-case bound. In this paper we show how we can merge two RLBWTs into a run-length compressed eBWT in $O(r)$ space and $\tilde{O}(r+L)$ time, where $r$ is the number of runs in the final eBWT and $L$ is the sum of the longest common prefix (LCP) values at the beginnings of those runs.
[…]
2 Preliminaries
For the sake of brevity, we assume readers are familiar with run-length compression, the Burrows-Wheeler Transform (BWT), run-length compressed BWTs (RLBWTs), suffix arrays (SAs) and longest common prefix (LCP) arrays. The extended BWT (eBWT) of two strings $S[1..m]$ and $T[1..n]$ is an interleaving $\mathrm{BWT}_{S,T}[1..m+n]$ of the characters in $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..m]$. For $1\leq h\leq m+n$, $\mathrm{BWT}_{S,T}[h]$ has the lexicographically $h$th largest context. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ are subsequences of $\mathrm{BWT}_{S,T}$, if it has $r$ runs then they each have at most $r$ runs.
The context of $\mathrm{BWT}_{S}[i]$ is
$\mathrm{context}_{S}(i)=S[\mathrm{SA}_{S}[i]..m]\circ S[1..\mathrm{SA}_{S}[i]-1]$
(unless $\mathrm{SA}_{S}[i]=1$, in which case the context is just $S[1..m]$) and the context of $\mathrm{BWT}_{T}[j]$ is
$\mathrm{context}_{T}(j)=T[\mathrm{SA}_{T}[j]..n]\circ T[1..\mathrm{SA}_{T}[j]-1]$
(unless $\mathrm{SA}_{T}[j]=1$, in which case the context is just $T[1..n]$), where $\circ$ denotes concatenation. If we assume $S$ and $T$ are each terminated by an end-of-string symbol that occurs nowhere else in $S$ and $T$ then two characters in the same BWT cannot have the same context, and a character in $\mathrm{BWT}_{S}$ and a character in $\mathrm{BWT}_{T}$ have the same context if and only if $S=T$ and the characters have the same positions (in which case we can break ties arbitrarily without affecting $\mathrm{BWT}_{S,T}$).
Given the $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}T[1..n]$ run-length compressed and each with at most $r$ runs, in $\tilde{O}(r)$ time we can build two $O(r)$-space data structures with which, given $i$ and $j$, we can compare $\mathrm{context}_{S}(i)$ and $\mathrm{context}_{T}(j)$ character by character in $\mathrm{polylog}(m+n)$ time per character until we find a mismatch. For example, given $\mathrm{BWT}_{S}$ as a sequence of (character, length) pairs, in $O(r\log r)$ time we can
  1. 1.
    add to each run’s pair its starting position in $\mathrm{BWT}_{S}$, to obtain a triple;
  2. 2.
    stably sort the triples lexicographically by character;
  3. 3.
    add to each triple the prefix sum of the lengths in the triples now before it.
The prefix sum stored in a triple is the value of the $\mathrm{LF}$ function for $S$ applied to the starting position stored in that triple, and the starting position is the value of the $\psi$ function applied to the prefix sum. Therefore, given $i$, we can
  1. 1.
    use binary search to find $i$’s predecessor $\mathrm{pred}(i)$ among the prefix sums in $O(\log r)$ time;
  2. 2.
    take the character stored in the triple with prefix sum $\mathrm{pred}(i)$ as the first character in $\mathrm{context}_{S}(i)$;
  3. 3.
    take the starting position stored in the same triple as $\psi(\mathrm{pred}(i))$;
  4. 4.
    compute the position
    $\psi(i)=\psi(\mathrm{pred}(i))+(i-\mathrm{pred}(i))$
    in $\mathrm{BWT}_{S}$ of the character whose context is $\mathrm{context}_{S}(i)$ cyclically shifted 1 character to the left.
We can repeat this process with $\psi(i)$ to obtain the second character of $\mathrm{context}_{S}(i)$, and so on. We could store a sparse bitvector with 1s marking the prefix sums and use rank and select queries on it to find $\mathrm{pred}(i)$, but we leave such optimizations to the full version of this paper as they change the time complexity by at most polylogarithmic factors.
3 Algorithm
Suppose we are to merge the RLBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$ of strings $S[1..m]$ and $T[1..n]$ into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ of $S$ and $T$ together. The pseudocode for our algorithm is shown in Figure 1.
$1i\leftarrow$ 1
[…]
3:while $i\leq m$ and $j\leq n$ do
4:  $k\leftarrow\mbox{length of leading run of $\mathrm{BWT}_{S}[i..m]$}$
5:  $\ell\leftarrow\mbox{length of leading run of $\mathrm{BWT}_{T}[j..n]$}$
6:  if $\mathrm{BWT}_{S}[i]=\mathrm{BWT}_{T}[j]$ then
7:   if $i+k\leq m$ and $\mathrm{context}_{S}(i+k)\prec\mathrm{context}_{T}(j+\ell-1)$ then
8:     binary search for the smallest value $j^{\prime}$ with $\mathrm{context}_{T}(j^{\prime})\succ\mathrm{context}_{S}(i+k)$
9:     output $\mathrm{BWT}_{S}[i..i+k-1]\circ\mathrm{BWT}_{T}[j..j^{\prime}-1]$
10:     $i\leftarrow i+k$
11:     $j\leftarrow j^{\prime}$
12:   else if $j+\ell\leq n$ and $\mathrm{context}_{T}(j+\ell)\prec\mathrm{context}_{S}(i+k-1)$ then
13:     binary search for the smallest value $i^{\prime}$ with $\mathrm{context}_{S}(i^{\prime})\succ\mathrm{context}_{T}(j+\ell)$
14:     output $\mathrm{BWT}_{S}[i..i^{\prime}-1]\circ\mathrm{BWT}_{T}[j..j+\ell-1]$
15:     $i\leftarrow i^{\prime}$
16:     $j\leftarrow j+\ell$
17:   else
18:     output $\mathrm{BWT}_{S}[i..i+k-1]\circ\mathrm{BWT}_{T}[j..j+\ell-1]$
19:     $i\leftarrow i+k$
20:     $j\leftarrow j+\ell$
21:   end if
22:  else
23:   if $\mathrm{context}_{S}(i)\prec\mathrm{context}_{T}(j)$ then
24:     if $\mathrm{context}_{S}(i+k-1)\prec\mathrm{context}_{T}(j)$ then
25:      output $\mathrm{BWT}_{S}[i..i+k-1]$
26:      $i\leftarrow i+k$
27:     else
28:      binary search for the smallest value $i^{\prime}$ with $\mathrm{context}_{S}(i^{\prime})\succ\mathrm{context}_{T}(j)$
29:      output $\mathrm{BWT}_{S}[i..i^{\prime}-1]$
30:      $i\leftarrow i^{\prime}$
31:     end if
32:   else
33:     if $\mathrm{context}_{T}(j+\ell-1)\prec\mathrm{context}_{S}(i)$ then
34:      output $\mathrm{BWT}_{T}[j..j+\ell-1]$
35:      $j\leftarrow j+\ell$
36:     else
37:      binary search for the smallest value $j^{\prime}$ with $\mathrm{context}_{T}(j^{\prime})\succ\mathrm{context}_{S}(i)$
38:      output $\mathrm{BWT}_{T}[j..j^{\prime}-1]$
39:      $j\leftarrow j^{\prime}$
40:     end if
41:   end if
42:  end if
43:end while
44:if $i\leq m$ then
45:  output $\mathrm{BWT}_{S}[i..m]$
46:else if $j\leq n$ then
47:  output $\mathrm{BWT}_{T}[j..n]$
48:end if
Figure 1: Pseudocode for merging $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$ into $\mathrm{BWT}_{S,T}[1..m+n]$.
Assume we have already merged $\mathrm{BWT}_{S}[1..i-1]$ and $\mathrm{BWT}_{T}[1..j-1]$ into $\mathrm{BWT}_{S,T}[1..i+j-2]$ correctly. When we first reach the while loop (lines 3–43) $i=j=1$ so this assumption is trivially true. If $i>m$ then $\mathrm{BWT}_{S,T}[i+j-1..m+n]=\mathrm{BWT}_{T}[j..n]$ (line 45) and if $j>n$ then $\mathrm{BWT}_{S,T}[i+j-1..m+n]=\mathrm{BWT}_{S}[i..m]$ (line 47), so assume $i\leq m$ and $j\leq m$ and consider only the while loop. Let $k$ and $\ell$ be the lengths of the leading runs of $\mathrm{BWT}_{S}[i..m]$ and $\mathrm{BWT}_{T}[j..n]$ respectively (lines 4 and 5).
First suppose $\mathrm{BWT}_{S}[i]=\mathrm{BWT}_{T}[j]$ (lines 7–21), meaning the leading runs $\mathrm{BWT}_{S}[i..i+k-1]$ and $\mathrm{BWT}_{T}[j..j+\ell-1]$ of $\mathrm{BWT}_{S}[i..m]$ and $\mathrm{BWT}_{T}[j..n]$ consist of copies of the same character. If there exists a character $\mathrm{BWT}_{S}[i+k]$ after the run $\mathrm{BWT}_{S}[i..i+k-1]$ and $\mathrm{context}_{S}(i+k)\prec\mathrm{context}_{T}(j+\ell-1)$ (lines 8–11) then $\mathrm{BWT}_{S}[i+k]$ precedes some non-empty suffix $\mathrm{BWT}_{T}[j^{\prime}..j+\ell-1]$ of the run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$, with $j^{\prime}$ being the smallest value with $\mathrm{context}_{T}(j^{\prime})\succ\mathrm{context}_{S}(i+k)$. We use binary search to find $j^{\prime}$ in the range $j,j+\ell)$ (line 8), output
$\mathrm{BWT}_{S}[i..i+k-1]\circ\mathrm{BWT}_{T}[j..j^{\prime}-1]=(\mathrm{BWT}_{S}[i])^{k+j^{\prime}-j}$
(line 9) and reset $i$ to $i+k$ (line 10) and $j$ to $j^{\prime}$ (line 11). Since the characters we output are all equal, we need not consider their order. The case when there exists a character $\mathrm{BWT}_{T}[j+\ell]$ after the run $\mathrm{BWT}_{T}[j..j+\ell-1]$ and $\mathrm{context}_{T}(j+\ell)\prec\mathrm{context}_{S}(i+k-1)$ (lines 13–16) is symmetric.
By the definition of the eBWT, it is impossible for $\mathrm{BWT}_{S}[i+k]$ to precede a non-empty suffix of the run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$ and for $\mathrm{BWT}_{T}[j+\ell]$ to precede a non-empty suffix of the run $\mathrm{BWT}_{S}[i..i+k-1]$ in $\mathrm{BWT}_{S,T}$ in $\mathrm{BWT}_{S,T}$, since then $\mathrm{BWT}_{S}[i+k]$ would have to both precede and follow $\mathrm{BWT}_{T}[j+\ell]$. Therefore, the only remaining case with $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$ is when either $i+k=m+1$ or $\mathrm{BWT}_{S}[i+k]$ follows the entire run $\mathrm{BWT}_{T}[j..j+\ell-1]$ in $\mathrm{BWT}_{S,T}$ and either $j+\ell=n+1$ or $\mathrm{BWT}_{T}[j]$ follows the entire run $\mathrm{BWT}_{S}[i..i+k-1]$ in $\mathrm{BWT}_{S,T}$. In this case, we output
$\mathrm{BWT}_{S}[i..i+k-1]\circ\mathrm{BWT}_{T}[j..j+\ell-1]=(\mathrm{BWT}_{S}[i])^{k+j}$
(line 18) and reset $i$ to $i+k$ (line 19) and $j$ to $j+\ell$ (line 20).
Now suppose $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$ (lines 23–42). Because $S[\mathrm{SA}_{S}[i]-1]=\mathrm{BWT}_{S}[i]$ (or $S[m]=\mathrm{BWT}_{S}[i]$ if $\mathrm{SA}_{S}[i]=1$ so $S[1..\mathrm{SA}_{S}[i]-1]$ is empty) and $T[\mathrm{SA}_{T}[j]-1]=\mathrm{BWT}_{T}[j]$ (or $T[n]=\mathrm{BWT}_{T}[j]$ if $\mathrm{SA}_{T}[j]=1$ so $T[1..\mathrm{SA}_{T}[j]-1]$ is empty), $\mathrm{context}_{S}(i)$ and $\mathrm{context}_{T}(j)$ differ on their last characters. Therefore, we consider only two cases: when $\mathrm{context}_{S}(i)\prec\mathrm{context}_{T}(j)$ (lines 24–31) and when $\mathrm{context}_{S}(i)\succ\mathrm{context}_{T}(j)$ (lines 33–40).
Suppose $\mathrm{context}_{S}(i)\prec\mathrm{context}_{T}(j)$, meaning $\mathrm{BWT}_{S}[i]$ precedes $\mathrm{BWT}_{T}[j]$ in $\mathrm{BWT}_{S,T}$. If $\mathrm{context}_{S}(i+k-1)\prec\mathrm{context}_{T}(j)$ then the entire run $\mathrm{BWT}_{S}[i..i+k-1]$ precedes $\mathrm{BWT}_{T}[j]$ in $\mathrm{BWT}_{S,T}$, so we output that run (line 25) and reset $i$ to $i+k$ (line 26). Otherwise, some non-empty prefix $\mathrm{BWT}_{S}[i..i^{\prime}-1]$ of the run $\mathrm{BWT}_{S}[i..i+k-1]$ precedes $\mathrm{BWT}_{T}[j]$ in $\mathrm{BWT}_{S,T}$ and some non-empty suffix $\mathrm{BWT}_{S}[i^{\prime}..i+k-1]$ follows it, with $i^{\prime}$ being the smallest value with $\mathrm{context}_{S}(i^{\prime})\succ\mathrm{context}_{T}(j)$. We use binary search (line 28) to find $i^{\prime}$ in the range (i,i+k1](i,i+k-1], comparing the context of a character in $\mathrm{BWT}_{S}[i+1..i+k-1]$ to $\mathrm{context}_{T}(j)$ at each step. We output $\mathrm{BWT}_{S}[i..i^{\prime}-1]$ (line 29) and reset $i$ to $i^{\prime}$ (line 30). The case when $\mathrm{context}_{S}(i)\succ\mathrm{context}_{T}(j)$ is symmetric.
4 Analysis
Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ both consist of at most $r$ runs, outputting $\mathrm{BWT}_{S}[i..m]$ (line 45) or $\mathrm{BWT}_{T}[j..n]$ (line 47) in run-length compressed form takes $O(r)$ time, where $r$ is the number of runs in $\mathrm{BWT}_{S,T}$, and we can focus on the complexity of the while loop.
It is not difficult to check that during each pass through the while loop, we output an entire run in $\mathrm{BWT}_{S,T}$. Whenever we reach line 9 we have $\mathrm{context}_{S}(i+k)\prec\mathrm{context}_{T}(j^{\prime})$ so, after we output
$\mathrm{BWT}_{S}[i..i+k-1]\circ\mathrm{BWT}_{T}[j..j^{\prime}-1]=(\mathrm{BWT}_{S}[i])^{k+j^{\prime}-j}\,$,
the next character we output is $\mathrm{BWT}_{S}[i+k]\neq\mathrm{BWT}_{S}[i]$ and it starts a new run in $\mathrm{BWT}_{S,T}$. Symmetrically, after we reach line 14 and output
$\mathrm{BWT}_{S}[i..i^{\prime}-1]\circ\mathrm{BWT}_{T}[j..j+\ell-1]=(\mathrm{BWT}_{S}[i])^{i^{\prime}-i+\ell}\,$,
the next character we output is $\mathrm{BWT}_{T}[j+\ell]\neq\mathrm{BWT}_{T}[j]$ and it starts a new run in $\mathrm{BWT}_{S,T}$. After we reach line 18 and output
$\mathrm{BWT}_{S}[i..i+k-1]\circ\mathrm{BWT}_{T}[j..j+\ell-1]=(\mathrm{BWT}_{S}[i])^{k+\ell}\,$,
the next character we output (if there is one) is either $\mathrm{BWT}_{S}[i+k]\neq\mathrm{BWT}_{S}[i]$ or $\mathrm{BWT}_{T}[j+\ell]\neq\mathrm{BWT}_{T}[j]$ and it starts a new run in $\mathrm{BWT}_{S,T}$. After we reach line 25 and output
$\mathrm{BWT}_{S}[i..i+k-1]=(\mathrm{BWT}_{S}[i])^{k}\,$,
the next character we output is either $\mathrm{BWT}_{S}[i+k]\neq\mathrm{BWT}_{S}[i]$ or $\mathrm{BWT}_{T}[j]\neq\mathrm{BWT}_{S}[i]$ and it starts a new run in $\mathrm{BWT}_{S,T}$. Symmetrically, after we reach line 34 and output
$\mathrm{BWT}_{T}[j..j+\ell-1]=(\mathrm{BWT}_{T}[j])^{\ell}\,$,
the next character we output is either $\mathrm{BWT}_{T}[i+\ell]\neq\mathrm{BWT}_{T}[j]$ or $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$ and it starts a new run in $\mathrm{BWT}_{S,T}$. After we reach line 29 and output
$\mathrm{BWT}_{S}[i..i^{\prime}-1]=(\mathrm{BWT}_{S}[i])^{i^{\prime}-i}\,$,
the next character we output is $\mathrm{BWT}_{T}[j]\neq\mathrm{BWT}_{S}[i]$ and it starts a new run in $\mathrm{BWT}_{S,T}$. Symmetrically, after we reach line 38 and output
$\mathrm{BWT}_{T}[j..j^{\prime}-1]=(\mathrm{BWT}_{T}[j])^{j^{\prime}-k}\,$,
the next character we output is $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$ and it starts a new run in $\mathrm{BWT}_{S,T}$. It follows that we make $O(r)$ passes through the while loop.
Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ and $\mathrm{BWT}_{S,T}$ are run-length compressed, the only operations that can take more than constant time during a single pass through the while loop are comparisons between characters’ contexts, of which we make $O(\log(m+n))$ during each pass. It is not difficult to check, however, that whenever we compare two characters’ contexts
  • those two characters are not equal, so they are in different runs in $\mathrm{BWT}_{S,T}$;
  • at least one of those two characters is in the run in $\mathrm{BWT}_{S,T}$ output during the current pass through the while loop or the run in $\mathrm{BWT}_{S,T}$ output during the next pass.
When we check whether $\mathrm{context}_{S}(i+k)\prec\mathrm{context}_{T}(j+\ell-1)$ in line 7,
$\mathrm{BWT}_{S}[i+k]\neq\mathrm{BWT}_{S}[i]=\mathrm{BWT}_{T}[j]=\mathrm{BWT}_{T}[j+\ell-1]\,.$
The comparison in line 12 is symmetric. When we perform the binary search in line 8, in each step we check whether $\mathrm{context}_{T}(j^{\prime})\succ\mathrm{context}_{S}(i+k)$ for some $j^{\prime}<j+\ell$ with
$\mathrm{BWT}_{T}[j^{\prime}]=\mathrm{BWT}_{T}[j]=\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{S}[i+k]\,.$
The binary search in line 13 is symmetric. When we check whether $\mathrm{context}_{S}(i)\prec\mathrm{context}_{T}(j)$ in line 23, $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$. When we check whether $\mathrm{context}_{S}(i+k-1)\prec\mathrm{context}_{T}(j)$ in line 24,
$\mathrm{BWT}_{S}[i+k-1]=\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]\,.$
The comparison in line 33 is symmetric. When we perform the binary search in line 28, in each step we check whether $\mathrm{context}_{S}(i^{\prime})\succ\mathrm{context}_{T}(j)$ for some $i^{\prime}<i+k$ with
$\mathrm{BWT}_{S}[i^{\prime}]=\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]\,.$
The binary search in line 37 is symmetric.
It follows that we can charge all the context comparisons to runs in $\mathrm{BWT}_{S,T}$ such that
  • each run in $\mathrm{BWT}_{S,T}$ has $O(\log(m+n))$ context comparisons charged to it;
  • if a comparison between two characters’ contexts is charged to a run in $\mathrm{BWT}_{S,T}$ then one of the characters is in that run and the other is not.
The number of characters we extract to compare the context of a character in a run $\mathrm{BWT}_{S,T}[p..q]$ to the context of a character in another run is at most
[…]
(unless $q=m+n$, in which case the number is just $\mathrm{LCP}_{S,T}[p]$), where $\mathrm{LCP}_{S,T}[1]=0$ and $\mathrm{LCP}_{S,T}[h]$ is the length of the longest common prefix of the $h$th and (h1)(h-1)st characters in $\mathrm{BWT}_{S,T}$. Therefore, the total time we spend comparing contexts is
[…]
If we set
[…]
then this time bound is $\tilde{O}(r+L)$. Notice we need not compute $L$ to run our algorithm.
Theorem 1
Given the run-length compressed eBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $\tilde{O}(r+L)$ time and $O(r)$ space, where $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of the values in $\mathrm{LCP}_{S,T}$ at the beginnings of runs in $\mathrm{BWT}_{S,T}$.